OpenStudy (anonymous):

slove the equation by subustion: x=y-5 and 3x+2y=3

OpenStudy (anonymous):

HELP!

OpenStudy (anonymous):

x = y - 5 3(y-5) + 2y = 3 3y - 15 + 2y = 3 5y = 18 y = 18/5 x = 18/5 - 5 x = -7/5 (-7/5, 18/5)